OpenStudy (anonymous):

5x-12=2x+6 sole equation?

OpenStudy (anonymous):

5x-2x = 6 + 12 3x = 18 x=6

OpenStudy (anonymous):

5x-12=2x+6 5x-2x = 6 + 12 3x = 18 x=6

OpenStudy (anonymous):

\[4t-7-t=7-2t-5\]

OpenStudy (anonymous):

p\[p+6+3p=1+3p +1\]

OpenStudy (anonymous):

\[\sqrt{54}\]

OpenStudy (anonymous):

find the coordinates of the midpoint of a segment with the given end points. j(-1,7) k(3,-3) the formula is \[x _{1}+x _{2}\div2 , y _{1}+y _{2} \div2\]

OpenStudy (ilovecake):

Do I get medal and fan Math Geek?